$SPAD/src/input as-eg6.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1995.

———-

— * —

\getchunk{license}
)co aseg6.as
)show C1
)show C2
)show C3
)show C4
)show D1
)show D2
)show D3
)show D4
)show D5
)show C2 Integer
)show C4 Integer
)show D2 Integer
)show D4 Integer
x: D5 Integer := 1
-- should be 3
x+x
-- should be 13
foof(12)$D3

———-

References^{[1] nothing}

4